On 10/04/18, consider a fixed-coupon bond whose features are the following: 

• face value: $1,000 

• coupon rate: 8% 

• coupon frequency: semiannual 

• maturity: 12/06/18 


What are the future cash flows delivered by this bond?
Answer #1

Considering Bond is issues on 10/04/2018


Future Cash flow will be

Face Value - $1000

Coupon Amount from (10/04/18 - 12/06/18 = 64 days) @8%p.a = 1000*8%*64/365 = $14.02


Total Future cash flow= $1014.02


answered by: anonymous